# Evolving Network Modeling Driven by the Degree Increase and Decrease Mechanism

### v1

@@ III-A Modeling of Evolving network
Service process. Any vertex can be connected to as many vertices as possible, indicating that the number of vertices under service is without limit, numerically from 00 to +$\infty$+ ∞. Therefore, the number of servers is infinite, and all servers work in parallel. Besides, the decrease of vertices is random, leading to a Poisson output flow. Hence, the service time under each server is exponentially distributed with a parameter $\mu$ which is the output rate, expressed as
@@
System capacity. Since there is no limit for the number of vertices in our proposed model, the vertex degree queueing system capacity is +$\infty$+ ∞.
@@ III-B Analyses for Degree Distribution
We analyze the degree of an arbitrary vertex $i$ in the network, then the degree distribution of vertex $i$ we deduced is the degree distribution of the network. The degrees of a vertex $i$ change with time, and the degree value can be taken as a stochastic process, denoted by $\{K(t),\ t\geq 0\}$. Furthermore, we assume that the future state of the degree value $K(t+\Delta t)$, given the past states and the present state $K(t)$, is independent of the past states and depends only on the present state, which is intuitive according to the increase and decrease mechanism for the vertex degree. Thus, the degrees of an arbitrary vertex $i$ varying with time denoted by $\{K(t),\ t\geq 0\}$ is known as a Markov chain whose state space is Ω={1, $\Omega=\{1,\ 2,\ \cdots\}$, and we let
@@ III-B1 Degree distribution with $\lambda(k)=k$
As the queueing system is stationary, according to the Kolmogorov backward equation, for $k$=1, 2, $\cdots$, we have
@@
Figure 2: An illustration of the state transition of a vertex (denoted by $i$)’s degree: As we can see above the arrow $f(k)$$(k=1,2,...)$ indicates the degree increase rate relevant to the degree $k$, and beneath the arrow, $k\mu$ indicates the degree decrease rate. The figures in circles indicate the degree value.
@@ III-B2 Degree distribution with $\lambda(k)=ln(1+k)$
We calculate $S(\mu)$ by numerical simulations. The sum should be calculated from $k=2$ to +$\infty$+ ∞ theoretically. However, in simulations, we set $k$ large enough, meanwhile the largest value of $k$ should avoid the underflow of items in $S(\mu)$. Then, set the value for $\mu$, by computer calculation we can obtain $S(\mu)$ numerically, and further get $P_{1}$ according to
@@ IV-A Modeling of Network
Primarily, we describe the construction of our network in simulations based on our model. In terms of time passing, the time is continuous, and the interval of events that new vertices come and existing vertices are deleted is generated by exponential distributions. For the connection, each new vertex carried with $m$ edges comes and connects to existing vertices in networks. Allocate new edges of newly coming vertices connecting to existing vertices according to the form of the degree increase rate $\lambda(k)$. For the first form $\lambda(k)$=$k$, we use the $\frac{k_{i}}{\sum^{j}k}$ as the probability for vertex $i$ connected to a new vertex. For the second form $\lambda(k)$=$ln(1+k)$, $\frac{ln(1+k_{i})}{\sum^{j}ln(1+k_{j})}$ is taken as the probability for vertex $i$ connected to a new vertex. We utilize $networkx$ packages in Python to carry out the simulation. The procedures of network construction are as follows:
@@
Based on the above algorithm, we construct networks demonstrated in Fig. 3. The sub-figures (a)-(c) show network snapshots with $\lambda(k)=k$, given $\mu$=1.01, 1.21, 1.41 respectively, and sub-figures (d)-(f) present network snapshots with $\lambda(k)=ln(1+k)$, given $\mu$=0.005, 0.01, 0.05. The initial number of vertices is set as $m_{0}$=10, and the edges of new vertices are $m$=3 for both forms of $\lambda(k)$. $T$ is set as 3.5 for networks in Fig. 3 (a)-(c) and 6.5 for (d)-(f), which is a reasonable length of time for network evolving in terms of the decrease rate $\mu$. The size and color of a vertex reflect its degree value. In terms of size, the bigger the size of the circle is, the larger degree it has. In terms of color, yellow indicates a large degree, while the other purple indicates a small degree, blue and green indicate a medium degree. The measure of color presenting the degree is only within the same network. As is shown in Fig. 3, networks in the sub-figures are all heterogenous, only a few vertices have a large degree value and the majority of vertices have small degree values. For both situations for the increase rate $\lambda(k)$ ($\lambda(k)$=$k$ and $\lambda(k)$=$ln(1+k)$), the network is getting sparser with the decrease rate $\mu$ increasing according to the snapshots of networks at the same time.
@@ IV-C1 Co-authorship network
As is illustrated in Fig. 8 marked by blue circles, the degree distribution of the co-authorship network is heterogenous The fraction is below 0.1 when $k$=4, and it is very small and stays steady when $k>7$ featured with a “long tail”. However, the degree distribution of the co-authorship network is not a typical Power-law distribution, as we can see in sub-figures (b) and (d). The degree distribution manifests as a curve going downward but not a straight line plot in the logarithmic coordinate. Figs. 8 (a) and (b) show the fitting results of $\lambda(k)=k$ where we set the parameter $\mu=1.28$ for the situation of $\lambda(k)=k$, and in Figs. 8 (c) and (d), we set the parameter $\mu=0.58$ for $\lambda(k)=ln(1+k)$. In both sub-figures, most circles are around the theoretical lines, suggesting that the theoretical distribution is in accordance with the practical degree distribution. Under the logarithm coordinate in sub-figures 8 (b) and (d), blue circles are not around the theoretical line when $k>10$. This is because the fraction of vertices with a degree larger than 10 is the same in the practical network owing to a small volume of data. Nevertheless, our theoretical distributions of both expressions for $ln(1+k)$ have a good fitting on the whole.
@@ IV-C2 Email communication network
In Figs. 9 (a) and (c), the practical distribution of the email communication network is also heterogenous. The degree distribution of this practical network is also not a Power-law distribution but a downward curve as we can see in the logarithm coordinate shown in Fig. 9 (b) and (d), yet it is featured with a “long tail” shown in (a) and (c). We set the parameter $\mu=1.05$ for the situation of $\lambda(k)=k$ presented in Figs. 9 (a) and (b), and set the parameter $\mu=0.40$ for $\lambda(k)=ln(1+k)$ illustrated in Figs. 9 (c) and (d). From Figs. 9 (a) and (c), we can see that the probability of $k$=1 is 0.3 according to the theoretical distribution of $\lambda(k)=k$, higher than the real degree distribution with a probability $P(k)$=0.2, however, they are close on the whole. Comparing sub-figures (a) and (c), the theoretical line with $\lambda=k$ is closer to real data circles than the line with $\lambda=ln(1+k)$, which suggests that the theoretical distribution with $\lambda(k)=ln(1+k)$ has better goodness of fitting than that with $\lambda(k)=ln(1+k)$ from the normal coordinate. From the plots under the logarithm coordinate shown in Fig. 10 (b), the theoretical line is like a straight line with a small curvature, indicating that the theoretical distribution is close to a Power-law distribution. The circle plot representing the practical degree distribution has a larger curvature compared to the theoretical line, which leads to the tail of the line does not agree with the circle plot. The situation in Fig. 9 (d) is different, the curvature of the theoretical line is larger than that of the practical distribution.
@@ IV-C3 Autonomous systems traffic-flow network
As is illustrated in Fig. 10, the degree distribution of the AS network is also a “long-tail” distribution, where the fraction of vertices with degree $k$=1 is over 0.25, and the fractions of $k$¿5 are all below 0.05. We set $\mu$=1.04 for the theoretical distribution with $\lambda(k)=k$, and $\mu$=0.44 for that with $\lambda(k)=ln(1+k)$, respectively demonstrated in Figs. 10 (a) and (c) in normal coordinates, (b) and (d) in logarithmic coordinates. For $\lambda(k)=k$, most circles are around the theoretical line under normal coordinates as is illustrated in Fig. 10 (a), suggesting that the theoretical distribution with $\lambda(k)=k$ fits the real degree distribution well. While under the logarithm coordinate in Fig. 10 (b), the trend of tail circles does not correspond to the theoretical line. This is because of the non-uniform scale of axes of logarithm coordinates which magnifies the difference between the real distribution and the theoretical one when $k$ is getting large. For $\lambda(k)=ln(1+k)$, under the normal coordinate, most circles are also around the theoretical line except for $k$=2 whose theoretical probability is 0.25 approximately, less than the practical fraction which is about 0.28. However, under the logarithm coordinate, the tail of the theoretical line with $\lambda(k)=ln(1+k)$ is too much below the blue circles Fig.10 (d), which shows a poor fitting result. In general, there is superb goodness of fit of theoretical distribution with $\lambda(k)=k$ compared to $\lambda(k)=ln(1+k)$ for the real degree distributions of AS network.

### v2

@@ III-A Modeling of Evolving network
Service process. Any vertex can be connected to as many vertices as possible, indicating that the number of vertices under service is without limit, numerically from 0 to $\infty$. Therefore, the number of servers is infinite, and all servers work in parallel. Besides, the decrease of vertices is random, leading to a Poisson output flow. Hence, the service time under each server is exponentially distributed with a parameter $\mu$ which is the output rate, expressed as
@@
System capacity. Since there is no limit for the number of vertices in our proposed model, the vertex degree queueing system capacity is $\infty$.
@@ III-B Analyses for Degree Distribution
We analyze the degree of an arbitrary vertex $i$ in the network, then the degree distribution of vertex $i$ we deduced is the degree distribution of the network. The degrees of a vertex $i$ change with time, and the degree value can be taken as a stochastic process, denoted by $\{K(t),\ t\geq 0\}$. Furthermore, we assume that the future state of the degree value $K(t+\Delta t)$, given the past states and the present state $K(t)$, is independent of the past states and depends only on the present state, which is intuitive according to the increase and decrease mechanism for the vertex degree. Thus, the degrees of an arbitrary vertex $i$ varying with time denoted by $\{K(t),\ t\geq 0\}$ is known as a Markov chain whose state space is $\Omega=\{1,\ 2,\ \cdots\}$, and we let
@@ III-B1 Degree distribution with $\lambda(k)=k$
As the queueing system is stationary, according to the Kolmogorov backward equation, for kk=1, 2, $\cdots$, we have
@@
Figure 2: An illustration of the state transition of a vertex (denoted by ii)’s degree: As we can see above the arrow f(k)f(k)(k=1,2,)(k=1,2,...) indicates the degree increase rate relevant to the degree $k$, and beneath the arrow, $k\mu$ indicates the degree decrease rate. The figures in circles indicate the degree value.
@@ III-B2 Degree distribution with $\lambda(k)=ln(1+k)$
We calculate $S(\mu)$ by numerical simulations. The sum should be calculated from $k=2$ to $\infty$ theoretically. However, in simulations, we set $k$ large enough, meanwhile the largest value of $k$ should avoid the underflow of items in $S(\mu)$. Then, set the value for $\mu$, by computer calculation we can obtain $S(\mu)$ numerically, and further get $P_{1}$ according to
@@ IV-A Modeling of Network
Primarily, we describe the construction of our network in simulations based on our model. In terms of time passing, the time is continuous, and the interval of events that new vertices come and existing vertices are deleted is generated by exponential distributions. For the connection, each new vertex carried with $m$ edges comes and connects to existing vertices in networks. Allocate new edges of newly coming vertices connecting to existing vertices according to the form of the degree increase rate $\lambda(k)$. For the first form $\lambda(k)$=$k$, we use the $\frac{k_{i}}{\sum^{j}k}$ as the probability for vertex $i$ connected to a new vertex. For the second form $\lambda(k)$=$ln(1+k)$, $\frac{ln(1+k_{i})}{\sum^{j}ln(1+k_{j})}$ is taken as the probability for vertex $i$ connected to a new vertex. We utilize networkxnetworkx packages in Python to carry out the simulation. The procedures of network construction are as follows:
@@
Based on the above algorithm, we construct networks demonstrated in Fig. 3. The sub-figures (a)-(c) show network snapshots with $\lambda(k)=k$, given $\mu$=1.01, 1.21, 1.41 respectively, and sub-figures (d)-(f) present network snapshots with $\lambda(k)=ln(1+k)$, given $\mu$=0.005, 0.01, 0.05. The initial number of vertices is set as $m_{0}$=10, and the edges of new vertices are mm=3 for both forms of $\lambda(k)$. $T$ is set as 3.5 for networks in Fig. 3 (a)-(c) and 6.5 for (d)-(f), which is a reasonable length of time for network evolving in terms of the decrease rate $\mu$. The size and color of a vertex reflect its degree value. In terms of size, the bigger the size of the circle is, the larger degree it has. In terms of color, yellow indicates a large degree, while the other purple indicates a small degree, blue and green indicate a medium degree. The measure of color presenting the degree is only within the same network. As is shown in Fig. 3, networks in the sub-figures are all heterogenous, only a few vertices have a large degree value and the majority of vertices have small degree values. For both situations for the increase rate $\lambda(k)$ ($\lambda(k)$=$k$ and $\lambda(k)$=$ln(1+k)$), the network is getting sparser with the decrease rate $\mu$ increasing according to the snapshots of networks at the same time.
@@ IV-C1 Co-authorship network
As is illustrated in Fig. 8 marked by blue circles, the degree distribution of the co-authorship network is heterogenous The fraction is below 0.1 when kk=4, and it is very small and stays steady when $k>7$ featured with a “long tail”. However, the degree distribution of the co-authorship network is not a typical Power-law distribution, as we can see in sub-figures (b) and (d). The degree distribution manifests as a curve going downward but not a straight line plot in the logarithmic coordinate. Figs. 8 (a) and (b) show the fitting results of $\lambda(k)=k$ where we set the parameter $\mu=1.28$ for the situation of $\lambda(k)=k$, and in Figs. 8 (c) and (d), we set the parameter $\mu=0.58$ for $\lambda(k)=ln(1+k)$. In both sub-figures, most circles are around the theoretical lines, suggesting that the theoretical distribution is in accordance with the practical degree distribution. Under the logarithm coordinate in sub-figures 8 (b) and (d), blue circles are not around the theoretical line when $k>10$. This is because the fraction of vertices with a degree larger than 10 is the same in the practical network owing to a small volume of data. Nevertheless, our theoretical distributions of both expressions for $ln(1+k)$ have a good fitting on the whole.
@@ IV-C2 Email communication network
In Figs. 9 (a) and (c), the practical distribution of the email communication network is also heterogenous. The degree distribution of this practical network is also not a Power-law distribution but a downward curve as we can see in the logarithm coordinate shown in Fig. 9 (b) and (d), yet it is featured with a “long tail” shown in (a) and (c). We set the parameter $\mu=1.05$ for the situation of $\lambda(k)=k$ presented in Figs. 9 (a) and (b), and set the parameter $\mu=0.40$ for $\lambda(k)=ln(1+k)$ illustrated in Figs. 9 (c) and (d). From Figs. 9 (a) and (c), we can see that the probability of kk=1 is 0.3 according to the theoretical distribution of $\lambda(k)=k$, higher than the real degree distribution with a probability P(k)P(k)=0.2, however, they are close on the whole. Comparing sub-figures (a) and (c), the theoretical line with $\lambda=k$ is closer to real data circles than the line with $\lambda=ln(1+k)$, which suggests that the theoretical distribution with $\lambda(k)=ln(1+k)$ has better goodness of fitting than that with $\lambda(k)=ln(1+k)$ from the normal coordinate. From the plots under the logarithm coordinate shown in Fig. 10 (b), the theoretical line is like a straight line with a small curvature, indicating that the theoretical distribution is close to a Power-law distribution. The circle plot representing the practical degree distribution has a larger curvature compared to the theoretical line, which leads to the tail of the line does not agree with the circle plot. The situation in Fig. 9 (d) is different, the curvature of the theoretical line is larger than that of the practical distribution.
@@ IV-C3 Autonomous systems traffic-flow network
As is illustrated in Fig. 10, the degree distribution of the AS network is also a “long-tail” distribution, where the fraction of vertices with degree kk=1 is over 0.25, and the fractions of kk¿5 are all below 0.05. We set $\mu$=1.04 for the theoretical distribution with $\lambda(k)=k$, and $\mu$=0.44 for that with $\lambda(k)=ln(1+k)$, respectively demonstrated in Figs. 10 (a) and (c) in normal coordinates, (b) and (d) in logarithmic coordinates. For $\lambda(k)=k$, most circles are around the theoretical line under normal coordinates as is illustrated in Fig. 10 (a), suggesting that the theoretical distribution with $\lambda(k)=k$ fits the real degree distribution well. While under the logarithm coordinate in Fig. 10 (b), the trend of tail circles does not correspond to the theoretical line. This is because of the non-uniform scale of axes of logarithm coordinates which magnifies the difference between the real distribution and the theoretical one when $k$ is getting large. For $\lambda(k)=ln(1+k)$, under the normal coordinate, most circles are also around the theoretical line except for kk=2 whose theoretical probability is 0.25 approximately, less than the practical fraction which is about 0.28. However, under the logarithm coordinate, the tail of the theoretical line with $\lambda(k)=ln(1+k)$ is too much below the blue circles Fig.10 (d), which shows a poor fitting result. In general, there is superb goodness of fit of theoretical distribution with $\lambda(k)=k$ compared to $\lambda(k)=ln(1+k)$ for the real degree distributions of AS network.